OpenStudy (anonymous):

Which ordered pairs are solutions to the inequality –3x + y ≥ 7? (–2, –3) (–1, 8) (0, 7) (1, 9) (3, 18)

OpenStudy (anonymous):

and this one Jude is placing boxes on a shelf that holds up to 80 lb. He has x small boxes that weigh 5 lb each and y large boxes that weigh 12 lb each. Jude needs at least 4 small boxes on the shelf. Which inequalities model the situation? Choose exactly three answers that are correct. 5x + 12y < 80 5x + 12y ≤ 80 x > 4 x ≥ 4 y > 0 y ≥ 0

OpenStudy (anonymous):

any idea

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

2nd 3rd and 5th for Q 1

OpenStudy (gorv):

small box weigth=5*x=5x weight of large box= 12*y=12y

OpenStudy (gorv):

total must be less than equal to 80 5x+12y<=80 and x>=4